# A cross-platform analysis of polarization and echo chambers in climate change discussions

### v2

@@ III Youtube analysis
We have analysed the dataset created from the YouTube videos quoted and referenced on Twitter. The information gathered for each post includes the transcription and description of each video, the corresponding channel, and the user comments. The comment information allows us to create a bipartite network between users and posts since we know the users who have commented on each video. We can project the bipartite network into a post network or a user network. In the post network, the weight corresponds to the number of users in common and, in the user network, to the posts in common. The graphs are undirected by construction in both cases. In Fig. S12, we display the network between YouTube posts for each dataset. We performed a community analysis using the greedy optimisation of modularity on the weighted network [75]. The organisation of communities is complex and fuzzy with several interconnections except for the Doñana network, where communities 00 and $1$ have a clear separation. By inspecting the videos with a large weighted degree, we have identified the communities with mainly climate change sceptics: community $2$ in COP27, communities $1$ and $2$ in IPCC, community $2$ in Climate Refugees, and community 00 in Doñana (Fig. S13).
Despite the communities being more separated, there is a strong interaction between them, indicating a certain level of engagement with opposed content. The exception is the Doñana network, where communities exhibit a clear separation, aligning with the pronounced polarization observed in the corresponding Twitter network. We computed polarization indices on the subnetworks using modularity and EI indices, similar to our approach for Twitter (Fig. 8). Only the COP27 and Doñana networks show higher levels of polarization compared to the others. Overall, despite confronting views on climate change, the interaction between posts remains robust. We conducted a sentiment and emotion analysis of the comments by the community (Fig. 9), revealing that negative comments dominate all posts. However, the posts related to sceptic content feature higher negative sentiments, together with more anger and disgust emotions. The sentiment and emotions of comments were calculated following [82]
Figure 8: Polarization in the Youtube dataset. Quantification of polarization in the post network through (a) modularity (b) E-I index and (c) adaptive E-I index. We have calculated the indices in the real networks and in 100 randomisations where we preserve the degree sequences.
Figure 9: Average sentiment of comments in YouTube posts by community. Average sentiment and emotions: (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana split by mainstream and sceptic communities.
We projected the bipartite graph into a user network to analyse the system from an individual interaction perspective. In Fig. S14, we plot the user interaction, where we focus on the comments of users within the posts of the sceptics and mainstream communities that we have previously separated. Each node corresponds to a user, and the colour to the community according to the greedy modularity optimisation. Despite the users being split into several communities, most of them fall into two to three large communities, which suggests a certain level of polarization. For instance, most users fall within the three largest communities for COP27.
We computed the polarization in the user network by splitting it into two parts using the METIS algorithm and calculating the values in the configuration graphs. Unlike in the post networks, we observe clear signs of polarization in all the networks (Fig. S15). Despite the strong interaction between posts with different views, it is clear that the comment patterns of users are highly polarized. If we aim to measure the polarization between individuals, the user network is the best approximation.
Figure 10: Distribution of user ideology in Youtube. Average ideology of the users split between mainstream and sceptic ideas in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. We assigned an ideology to each user according to the most common community of its comments.
We computed the ideology distribution of users by analysing the communities where they post comments. Focusing on the users with four or more comments, we calculated each user’s ideology as the average ideology of the videos they commented on, weighted by the size of each community to address any imbalances. We considered the mainstream community equal to $1$, and the sceptic community equal to 00. Both distributions are normalised separately. Overall, they point out that there is a clear separation in the ideology of users according to the posts where they have commented.
IV Cross-platform analysis
Finally, we conducted an analysis linking Twitter and YouTube social media platforms. We computed the frequency with which each Twitter community references YouTube posts and their corresponding communities. To address the heterogeneity in community sizes observed in the YouTube dataset, we normalised the number of references to each YouTube community by the community size. Fig. 11 illustrates a strong connection between Twitter and YouTube communities, as most references in each Twitter community align with a specific YouTube community. Additionally, each Twitter community demonstrates a unique pattern of references to YouTube.
In the COP27 network, the larger left-biased Twitter community predominantly references the YouTube community 00, while the larger right-biased community solely references posts in the YouTube community $1$. Conversely, in the IPCC network, the larger right-biased community references various YouTube communities, indicating that sceptic YouTube posts are dispersed across different communities. Except for the IPCC network, posts shared by the larger right-biased community on Twitter predominantly align with a single YouTube community. Furthermore, sceptic posts on YouTube receive minimal references from other Twitter communities.
@@ IV Cross-platform analysis
Figure 11: Interaction between Twitter and Youtube communities. Number of references to YouTube communities by the top 10 Twitter communities in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. To account for the uneven size of YouTube communities, we have divided the counts by the YouTube community size.
To inspect in detail the patterns of YouTube post sharing, we conducted a similar analysis to the echo chamber methodology developed for the Twitter datasets. Using the community assignment of posts between sceptics and mainstream, labelled as 00 and $1$ respectively, we computed the average ideology of users and their neighbourhoods. The density plot in Fig. 12 has clearly defined echo chambers, with two centres on opposite sides of the spectrum. These results suggest that users close to each other share similar content. However, there are differences across the typology of videos shared, as users who share sceptical videos also share mainstream videos to some extent.
Figure 12: Twitter echo chambers in the sharing of YouTube links. Density plots of the average ideology of the videos shared by the neighbours of a user as a function of its average ideology.
V Discussion
In this study, we have assessed the dynamics of polarization and echo chambers on climate change discussions around four different events. Our findings reveal that the extent and nature of polarization can vary significantly depending not only on the specific topic but also on contextual factors. Interestingly, we noted an over-representation of left-biased media in all datasets except one, indicating the dominance of left-leaning perspectives in climate change discussions, possibly affecting media attention from right-aligned outlets. In discussions more directly related to partisan political contexts, such as Doñana and COP27, we observed a stronger polarization and distinct echo chambers than the discussion focused on factual information, centred around the IPCC’s Report. This suggests that proximity to partisan politics amplifies polarization and the formation of echo chambers within online climate change discourse. This may indicate a dynamic of partisan sorting, as already found in the Finnish Twittersphere [77]. Our analysis of echo chambers revealed nuanced findings, particularly regarding interaction types. While the networks of user mentions in COP27 and IPCC discussions showed minimal evidence of echo chambers, retweet networks prominently featured echo chamber dynamics, indicating a higher level of endorsement associated with retweets compared to other interaction types. Expanding our analysis to include YouTube videos referenced in Twitter datasets provided a cross-platform perspective on climate change discourse. We identified unique communities within YouTube videos, separating sceptic perspectives from mainstream content. Interestingly, alignment between Twitter and YouTube communities is strong, with Twitter communities exhibiting distinct preferences for sharing YouTube content, especially for the sceptic community. Thus, indicating the presence of cross-platform echo chambers. Our work suggests that polarization is ubiquitous across social media platforms, although the typology of interactions impacts the overall polarization.
@@ V Discussion
The Twitter data are made available in accordance with Twitter’s terms of service at https://github.com/clint-project/clint-data and the codes to calculate the metrics at https://github.com/clint-project/polarization_metrics.
@@ Acknowledgements
Appendix A Dataset description
@@
  • COP27: COP27, climate, COP
  • IPCC: IPCC, report, climate.
  • Climate Refugees: Refugee, climate.
  • Doñana: Doñana, clima.
@@
Figure S2: Timeseries of comments in the Youtube datasets. Number of comments per date in the Youtube datasets of COP27, IPCC, Climate Refugees and Doñana. For visualization purposes, we show the comments after 2021.
Appendix B Influencers by community
We report in Tables S1, S2, S3 and S4 the top influencers by community for the COP27, IPCC, Climate Refugees, and Doñana datasets, respectively.
@@ Appendix B Influencers by community
Table S1: Top 5 most relevant users by total inflow in the ten largest communities of the COP27 dataset
@@
Table S2: Top 5 most relevant users by total inflow in the ten largest communities of the IPCC dataset
@@
Table S3: Top 5 most relevant users by total inflow in the ten largest communities of the Climate Refugees dataset
@@
Table S4: Top 5 most relevant users by total inflow in the ten largest communities of the Doñana dataset
@@
We report in Tables S5, S6, S7 and S8 the most used languages by community for the COP27, IPCC, Climate Refugees, and Doñana datasets, respectively.
@@
Appendix C Additional results polarization and entropy in Twitter
In Fig. S3 we report the standarized polarization in the COP27, IPCC, Climate Refugees and Doñana networks calculated following [22] as
$\hat{\Phi}_{z}(G)=\frac{\Phi(G)-\Phi(G_{CM})}{\sqrt{<\Phi(G_{CM})^{2}>-<\Phi(G_{CM})>^{2}}}.$ (S1)
@@ Appendix C Additional results polarization and entropy in Twitter
Figure S3: Standarized polarization in the Twitter dataset. Standardized measures of polarization in the retweet networks after the denoising and the calculation of z-scores.
In Fig. S4 we show the outflow entropy as a function of the outgoing isolation for the mention network between communities. We also show in Fig. S5 we show the inflow entropy as a function of the incoming isolation for the retweet network between communities.
Figure S4: Community connectivity and entropy in mention networks. Outflow entropy as a function of the normalized outgoing connectivity in the mention networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of attainable values.
Figure S5: Community connectivity and entropy in retweet networks. Inflow entropy as a function of the normalized incoming connectivity in the retweet networks of COP27, IPCC, Climate Refugees, and Doñana. The grey area corresponds to the range of attainable values.
@@
We have computed in Fig. S6 the polarization scores in the subgraph between the two largest communities with opposing bias that speak the same language.
@@
Figure S6: Standarized polarization in Twitter by interaction type. Standardized modularity and the E-I indices in (a) the retweet and (b) mention networks. We computed the polarization over the subnetwork of the two largest communities of opposed political bias. We have calculated the indices in the real network, in 100 realizations of the configuration model with preserved degree sequences, and standardized them.
In Fig. S7 we report the bias echo chambers in the mention network and in Fig. S8 the reliability echo chambers in the retweet network.
@@
Figure S7: Detection of echo chambers in the mention network for political bias. Political bias of the neighborhood as a function of the bias of a user in the mention network for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
@@
Figure S8: Detection of echo chambers in the retweet network for reliability. Media reliability of the neighborhood as a function of the bias of a user in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. studied by considering only retweets.
@@
The chamber overlap can distinguish between users sharing similar echo chambers (Fig.S9). We constructed the user-to-user matrices where the values correspond to their chamber overlap. Using a hierarchical agglomerative algorithm, we clustered the distance matrix and obtained an almost perfect separation of users by communities. The separation between the users in the communities is more neat in the COP27 and Doñana datasets compared to IPCC and Climate Refugees. Nevertheless, we observe a clear trend of similar echo chambers among users within the same community. In the COP27 network, the pattern is more evident for the users in the mainstream community (in blue). The separation between communities and the matching of echo chambers suggest a clearer separation in the Doñana dataset.

### v3

@@ II.4 Cross-platform analysis
We have analysed the dataset created from the YouTube videos quoted and referenced on Twitter to assess if the observed polarisation facets are aligned with content consumption in video platforms as partisan media affects the dynamics of social networks [79]. The information gathered for each post includes the transcription and description of each video, the corresponding channel, and the user comments. The comment information allows us to create a bipartite network between users and posts since we know the users who have commented on each video. We can project the bipartite network into a post network or a user network. In the post network, the weight corresponds to the number of users in common and, in the user network, to the posts in common. The graphs are undirected by construction in both cases. We performed a community analysis using the greedy modularity optimisation on the weighted network [69]. The organisation of communities do not display a clear separation between communities with several interconnections except for the Doñana network, where communities zero and one have a clear separation (see Section S4).
@@
Figure 7: Interaction between Twitter and YouTube communities. Number of references to YouTube communities by the top six Twitter communities in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. To account for the uneven size of YouTube communities, we have divided the counts by the YouTube community size.
We conducted an analysis linking Twitter and YouTube social media platforms to study the cross-platforms effects. We create the Twitter subnetwork of users who shared YouTube links related to climate change, tagged according to the YouTube community of the videos. The connection between network organisation and the shared YouTube content is clearer in the Doñana and IPCC networks. In both cases, users tweeting videos from YouTube community zero are more tightly connected than other users (see the subnetwork in Figs. S22 and S23).
Furthermore, we computed the frequency with which each Twitter community references YouTube posts and their corresponding communities. To account for the heterogeneity in community sizes observed in the YouTube dataset, we normalised the number of references to each YouTube community by its size. A strong connection exists between Twitter and YouTube communities, as most references within each Twitter community align with a specific YouTube community. Additionally, each Twitter community exhibits a unique pattern of references to YouTube.
In the COP27 network, the larger left-wing biased Twitter community 0 predominantly references YouTube community 0 ($35.8\%$), while the larger right-wing biased and low-reliability community 1 primarily references posts in YouTube community 2 ($57.3\%$). Similarly, in the Climate Refugees network, Twitter community 6, known for spreading low-reliability information, interacts significantly with YouTube community 2 ($74.6\%$). Finally, in the Doñana network, community 0 from both social networks interacts intensively ($84.3\%$). Conversely, in the IPCC network, the larger right-wing biased community 1 references multiple YouTube communities, specifically communities 1 ($20.2\%$), 2 ($35.3\%$), and 5 ($23.4\%$), suggesting a tendency to share diverse typologies of YouTube content. With the exception of the IPCC network, posts shared by the larger right-wing biased community on Twitter predominantly align with a single YouTube community (Fig. 7).
III Discussion
Our results confirm that polarisation in climate change discussions manifests in different forms through structural, ideological, and emotional components. We also observed distinct views on climate change and patterns of content consumption. The structural polarisation around the subject is aligned with distinct political biases and information reliability consumption. Not only that, the topological communities also show an alignment on perspective towards the environmental crisis and a stronger emotional response. Overall, polarisation takes multiple aligned forms. While previous work often emphasised the role of ideological segregation, our cross-platform analysis shows that inter-community exposure remains significant—especially via mentions—but is frequently accompanied by affective repulsion and asymmetric negativity, particularly from low-reliability, right-leaning communities.
This suggests that polarisation is not merely a consequence of echo chambers or algorithmic filtering, but also reflects active processes of antagonism and identity signalling, in line with theories of affective polarisation. Even in multilingual and transnational contexts, such as the COP27 dataset, we observe stable polarisation structures shaped more by political affiliations and emotional tone than by language or geography. Moreover, the structural signature of interactions varies across platforms and interaction types: while retweets act as endorsement signals and tend to reinforce in-group cohesion, mentions frequently bridge ideological divides, albeit often with antagonistic or negative sentiment.
The interplay between information reliability and political alignment also emerges as a key factor. Right-leaning communities consistently share lower-reliability content and exhibit higher isolation and negativity, yet they remain engaged with mainstream communities—albeit often in confrontational terms. This pattern challenges the notion that low-reliability information spreads in isolation and points instead to a contested media space where political identity, trust, and emotion are tightly intertwined. Our results align with other studies showing that, despite higher engagement within ideologically similar groups, users are still exposed to opposing views [17]. However, such exposure does not necessarily mitigate polarisation. On the contrary, it has been shown that confrontation with opposing views can intensify polarisation instead of reducing it [21]. This is reflected in our sentiment analysis, where messages from right-biased communities consistently show higher levels of negativity—especially when directed toward left-leaning users [60]. Overall, the sentiment in the low-reliability community was more negative. Our results are in line with recent research [26], which finds that climate change dissenters are closely related and tend to engage with a left-biased mainstream community [26, 60, 80]. Furthermore, toxicity levels are not evenly distributed across interaction types: replies and mentions are significantly more toxic than other forms of engagement [20]. This observation suggests that low-reliability individuals may reference the mainstream community to engage in discussions but get little attention in return [20], highlighting an asymmetry between left- and right-leaning communities [81].
The cross-platform alignment between Twitter and YouTube reinforces the idea that ideological communities are transversal across platforms, with users not only engaging with like-minded individuals but also selecting and amplifying content from ideologically aligned sources elsewhere. This calls for a broader, integrated approach to studying online polarisation, moving beyond single-platform analyses. Additionally, our results indicate that the degree and manifestation of polarisation can vary substantially across topics and events—suggesting that it is not static, but shaped by contextual factors such as the political salience of the issue or the type of event. This may reflect a dynamic of partisan sorting, whereby individuals increasingly align their social media behaviour with broader political identities, as already observed in the Finnish Twittersphere [71]. Prior studies, such as [54], observed that the 2019 IPCC Special Report on Climate Change and Land sparked significant public discourse on social media, with dietary recommendations becoming a highly polarised topic. Similar to our findings on localised polarisation patterns (e.g., Doñana), that report showed that climate solutions affecting lifestyle choices may trigger elevated levels of contention on platforms like Twitter. The alignment between platforms could be amplified by recommendation algorithms, which can function as filter bubbles that reinforce users’ beliefs [13, 82]. However, recent studies have downplayed the contribution of algorithms to polarisation [83]. Our results support the idea that polarisation is ubiquitous across platforms, but not driven solely by lack of exposure to opposing viewpoints.
These findings have important implications for the broader climate change debate. The persistence of polarised, emotionally charged interactions—even in the presence of exposure to opposing views—suggests that efforts to foster consensus around climate policy must account for the affective and identity-driven nature of public discourse. Moreover, the asymmetric structure of negativity and the association between political alignment and information reliability can contribute to public confusion, distrust in science, and a fragmentation of climate narratives, making it harder to establish the shared understanding necessary for effective action. Addressing polarisation in this context requires more than correcting low-reliability information—it demands strategies that bridge ideological divides and rebuild trust in both institutions and science.
Our static analysis of the polarisation around climate change could be extended to assess how it evolves with time as a function of internal dynamics, external sources, and the unfolding of events. For instance, coordinated campaigns have been observed to impact social media dynamics [84], and could play a role in shaping the emergence or intensification of polarisation. It is unclear whether message negativity is intrinsic to the topic or evolves with the discussion, potentially triggered by specific interactions. The asymmetric spread of emotionally polarised content suggests that by targeting specific small groups, polarisation could be significantly decreased [80]. However, the opacity of most online social networks makes content moderation a major challenge. It would be essential to assess whether content moderation or fact-checking could have a positive impact on the reduction of polarisation. Since our results suggest that individuals sharing low-reliability information are exposed to opposing views, the development of intervention strategies becomes a major challenge. The role of exposure in increasing polarisation could be further addressed by investigating the temporal evolution of user ideology and reliability as a function of their progressive interactions with others. In a similar fashion, it opens the door to the development of models to unveil how the exposure to opposing views can further exacerbate polarisation.
@@ III Discussion
The Twitter data are made available in accordance with Twitter’s terms of service at https://github.com/clint-project/clint-data (https//doi.org:/10.5281/zenodo.14095218) and the codes to calculate the metrics at https://github.com/clint-project/polarization_metrics (https://doi.org/10.5281/zenodo.14095108). The YouTube data is available at
https://doi.org:/10.5281/zenodo.15046657.
@@ Acknowledgements
Appendix A Dataset collection and description
@@
We also provide the detailed json queries used in the extraction in Listings LABEL:lst:cop27, LABEL:lst:ipcc, LABEL:lst:climate-refugees and LABEL:lst:donana.
Listing 1: COP27 Query Parameters
1{
2 "start-time": "2022-09-01",
3 "end-time": "2022-11-28",
4 "keywords": [
5 "COP27",
6 "TogetherForImplementation",
7 "COPTV",
8 "PRECOP",
9 "PRECOP27",
10 "LossAndDamage",
11 "Egypt_COP27",
12 "COP27Egypt",
13 "ClimateAction",
14 "ClimateCrisis",
15 "ClimateJustice"
16 ]
17}
Listing 2: IPCC Query Parameters
1{
2 "start-time": "2023-03-18",
3 "end-time": "2023-03-27",
4 "keywords": [
5 "ipcc",
6 "@ipcc_ch"
7 ]
8}
Listing 3: Climate Refugees Query Parameters
1[
2 {
3 "start-time": "2008-01-01",
4 "end-time": "2023-01-01",
5 "keywords": [
6 "climate (refugees OR refugee)",
7 "climate migration",
8 "climate displaced"
9 ]
10 }
11]
Listing 4: Doñana Query Parameters
1{
2 "start-time": "2019-01-01",
3 "end-time": "2023-05-01",
4 "keywords": [
5 "Do~nana"
6 ]
7}
The YouTube videos appearing in the Twitter dataset were filtered using specific keywords for each dataset. The keywords used were as follows: COP27 – ”COP27”, ”climate”, and ”COP”; IPCC – ”IPCC”, ”report”, and ”climate”; Climate Refugees – ”Refugee” and ”climate”; and Doñana – ”Doñana” and ”clima.”
@@
Figure S2: Timeseries of comments in the Youtube datasets. Number of comments per date in the Youtube datasets of COP27, IPCC, Climate Refugees and Doñana. For visualisation purposes, we show the comments after 2021.
Appendix B Network analysis and structural polarisation
S2 Influencers by community
In this section, we identify the top influencers within different communities for each of the key datasets: COP27, IPCC, Climate Refugees, and Doñana. The influencers are ranked by their total inflow of interactions, highlighting those users who have the most significant impact within their respective communities. Tables S1, S2, S3, and S4 present the top 4 most relevant users in the ten largest communities for each dataset.
@@ Appendix B Network analysis and structural polarisation
Table S1: Top 4 most relevant users by total inflow in the ten largest communities of the COP27 dataset
@@
Table S2: Top 4 most relevant users by total inflow in the ten largest communities of the IPCC dataset
@@
Table S3: Top 4 most relevant users by total inflow in the ten largest communities of the Climate Refugees dataset
@@
Table S4: Top 4 most relevant users by total inflow in the ten largest communities of the Doñana dataset
@@
We present the most frequently used languages across the largest communities for each dataset in Tables S5, S6, S7, and S8, which correspond to the COP27, IPCC, Climate Refugees, and Doñana datasets, respectively.
@@
The language distribution across the largest communities in the COP27, IPCC, Climate Refugees, and Doñana datasets reveals notable patterns. These patterns not only reflect the diversity of linguistic representation within each dataset but also provide insights into the global reach.
S2 Analysis of global structural polarisation
To assess the degree of polarisation within the online discourse surrounding key climate-related topics, we calculate the standardised polarisation in the COP27, IPCC, Climate Refugees, and Doñana networks, following the method outlined in [34]. The polarisation is standardised by comparing the observed polarisation $\Phi(G)$ against the expected polarisation $\Phi(G_{CM})$ from a corresponding null model, resulting in a z-score that accounts for network-specific variations. The formula for this calculation is given by:
$\hat{\Phi}_{z}(G)=\frac{\Phi(G)-\Phi(G_{CM})}{\sqrt{\langle\Phi(G_{CM})^{2}\rangle-\langle\Phi(G_{CM})\rangle^{2}}}.$ (S1)
Figure S3 presents the standardised polarisation across the retweet networks for the mentioned datasets, illustrating the extent of polarisation after denoising and z-score calculation.
@@
Figure S3: Standardised polarisation in the Twitter dataset. Standardized measures of polarisation in the retweet networks after the denoising and the calculation of z-scores.
@@
Furthermore, we have computed in Fig. S4 the polarisation scores in the subgraph between the two largest communities with opposing bias that speak the same language.
@@
Figure S4: Standarized polarisation in Twitter by interaction type. Standardized modularity and the E-I indices in (a) the retweet and (b) mention networks. We computed the polarisation over the subnetwork of the two largest communities of opposed political bias. We have calculated the indices in the real network, in 100 realizations of the configuration model with preserved degree sequences, and standardized them.
S2 Bias and reliability
Figure S5: Distribution of bias and reliability in the largest communities.(a-d) Distribution of bias and reliability in the seven largest communities for (a) COP27, (b) IPCC, (c) Climate Refugees and (d) Doñana.
We tested the significance of the observed bias (Fig. S6) and reliability (Fig. S7) to determine whether the differences identified in our analysis are statistically meaningful. In both cases, the differences between the two selected communities in each network are statistically significant.
Figure S6: Mann-Whitney U test between the bias of the top 7 communities. Result of the Mann-Whitney U test between the bias of the top 7 communities. The test was conducted to assess if the community in the vertical axis is larger than the community in the horizontal axes. Asterisks report the significance of the tests (*p-value<<0.05, **p-value<<0.01 and ***p-value<<0.001).
Figure S7: Mann-Whitney U test between the reliability of the top 7 communities. Result of the Mann-Whitney U test between the reliability of the top 7 communities. The test was conducted to assess if the community in the vertical axis is larger than the community in the horizontal axes. Asterisks report the significance of the tests (*p-value<<0.05, **p-value<<0.01 and ***p-value<<0.001).
S2 Isolation and entropy
To evaluate the isolation and interaction of communities, we calculated the directed interactions between them, and we computed the normalised outflow difference given by
$\Delta F^{out}_{i}=\frac{\sum_{j,j\neq i}T_{ij}-T_{ii}}{\sum_{\forall j}T_{ij}},$ (S2)
and the entropy of flows as
$H^{out}_{i}=-\sum_{\forall j}p_{ij}logp_{ij},$ (S3)
where $T_{ij}$ is the number of interactions from community $i$ to community $j$ and $p_{ij}$ is the probability of having a retweet from $i$ to $j$ given by $T_{ij}/\sum_{\forall j}T_{ij}$. While the first index quantifies the ratio between external and internal links and goes from 1-1 to $1$ when all flows are external, the entropy provides information on the variety of those flows. The entropy is 0 when there is no variety of flows and $1$ when it is maximum. In Fig. S8, we represent the outflow entropy as a function of the isolation for the four networks.
We computed the minimum and maximum attainable entropy values as a function of $\Delta F^{out}_{i}$ and $p_{ii}$ for orientation. Given a value of $\Delta F^{out}_{i}$ we have
$\frac{\sum_{j,j\neq i}T_{ij}}{{\sum_{\forall j}}T_{ji}}=\Delta F_{i}+\frac{T_{ii}}{{\sum_{\forall j}T_{ij}}}=\Delta F_{i}+p_{ii}.$ (S4)
The minimum value of the entropy corresponds to the situation when a community is connected only to itself and one external community (except for the extreme case $\Delta F_{i}=1$), which we can plug into the entropy as
$H^{in}_{i}=-p_{ii}logp_{ii}-\Delta F_{i}+p_{ii}log\Delta F_{i}+p_{ii}.$ (S5)
$H^{in}_{i}=-p_{ii}logp_{ii}-\sum^{N_{c}-1}_{j=0}\frac{\Delta F_{i}+p_{ii}}{N_{c}-1}log\frac{\Delta F_{i}+p_{ii}}{N_{c}-1}.$ (S6)
In Figs. S8-S9 we show the outflow entropy as a function of the outgoing isolation for the retweet and mention networks between communities.
Figure S8: Community connectivity and entropy in retweet networks. Outflow entropy as a function of the normalised outgoing isolation in the retweet networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of feasible values.
Figure S9: Community connectivity and entropy in mention networks. Outflow entropy as a function of the normalised outgoing connectivity in the mention networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of attainable values.
We also show in Fig. S10 the inflow entropy as a function of the incoming isolation for the retweet network between communities.
Figure S10: Community connectivity and entropy in retweet networks. Inflow entropy as a function of the normalized incoming connectivity in the retweet networks of (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. The grey area corresponds to the range of attainable values.
We report in Tables S9 and S10 we report the values of isolation and entropy for the 8 largest communities in each network.
Network Interaction Community $\Delta F^{out}$ $H^{out}$ Interaction Community $\Delta F^{out}$ $H^{out}$
0 -0.68 0.14 0 -0.57 0.17
1 -0.98 0.01 1 0.45 0.24
2 -0.72 0.10 2 0.03 0.15
IPCC Retweet 3 -0.61 0.12 Mention 3 0.73 0.19
4 -0.50 0.16 4 0.83 0.12
5 -0.76 0.09 5 0.26 0.24
6 -0.60 0.14 6 0.58 0.15
7 -0.43 0.18 7 0.51 0.18
0 -0.76 0.10 0 -0.50 0.17
1 -0.88 0.05 1 0.37 0.25
2 -0.97 0.01 2 0.18 0.19
COP27 Retweet 3 -0.94 0.03 Mention 3 -0.78 0.08
4 -0.60 0.10 4 0.09 0.22
5 -0.75 0.09 5 -0.32 0.18
6 -0.50 0.14 6 0.04 0.21
7 -0.51 0.13 7 0.01 0.20
0 -0.95 0.03 0 0.28 0.20
1 -0.48 0.16 1 -0.02 0.25
2 -0.33 0.17 2 0.55 0.27
Doñana Retweet 3 -0.33 0.18 Mention 3 0.47 0.29
4 -0.59 0.13 4 -0.31 0.19
5 -0.49 0.14 5 0.08 0.23
6 -0.32 0.19 6 0.70 0.24
7 -0.82 0.06 7 0.20 0.23
0 -0.67 0.10 0 -0.05 0.24
1 -0.58 0.13 1 0.24 0.27
2 -0.71 0.08 2 0.55 0.30
Climate Refugees Retweet 3 -0.68 0.10 Mention 3 -0.50 0.16
4 -0.57 0.12 4 0.33 0.26
5 -0.62 0.11 5 0.07 0.25
6 -0.93 0.03 6 -0.13 0.21
7 -0.72 0.08 7 -0.55 0.13
Table S9: Isolation and entropy metrics for the IPCC, COP27, Doñana, and Climate Refugees datasets, categorised by retweet and mention interactions for each community in terms of outflows.
Network Interaction Community $\Delta F^{in}$ $H^{in}$ Interaction Community $\Delta F^{in}$ $H^{in}$
0 -0.60 0.17 0 0.29 0.39
1 -0.97 0.02 1 -0.68 0.12
2 -0.73 0.12 2 -0.73 0.12
IPCC Retweet 3 -0.48 0.18 Mention 3 0.72 0.32
4 -0.35 0.21 4 0.60 0.33
5 -0.91 0.05 5 0.16 0.22
6 -0.63 0.14 6 0.74 0.35
7 -0.55 0.16 7 0.17 0.23
0 -0.69 0.12 0 -0.28 0.25
1 -0.83 0.07 1 0.06 0.22
2 -0.98 0.01 2 -0.68 0.08
COP27 Retweet 3 -0.93 0.03 Mention 3 -0.58 0.13
4 -0.69 0.09 4 -0.02 0.23
5 -0.82 0.07 5 0.25 0.29
6 -0.46 0.16 6 0.49 0.33
7 -0.54 0.14 7 0.11 0.25
0 -0.95 0.03 0 0.28 0.26
1 -0.41 0.17 1 -0.18 0.23
2 -0.48 0.15 2 0.02 0.24
Doñana Retweet 3 -0.36 0.18 Mention 3 0.40 0.29
4 -0.60 0.14 4 -0.31 0.21
5 -0.38 0.16 5 0.38 0.24
6 -0.49 0.15 6 0.41 0.28
7 -0.86 0.05 7 0.16 0.23
0 -0.65 0.11 0 0.15 0.27
1 -0.49 0.15 1 -0.02 0.26
2 -0.70 0.10 2 0.40 0.27
Climate Refugees Retweet 3 -0.67 0.11 Mention 3 -0.24 0.22
4 -0.52 0.14 4 0.70 0.32
5 -0.70 0.10 5 -0.10 0.23
6 -0.96 0.02 6 -0.21 0.18
7 -0.87 0.05 7 -0.71 0.09
Table S10: Isolation and entropy metrics for the IPCC, COP27, Doñana, and Climate Refugees datasets, categorised by retweet and mention interactions for each community in terms of inflows.
S2 Echo chambers
Echo chambers emerge in online discourse when users predominantly interact with like-minded individuals, reinforcing their existing beliefs. In political discussions, this phenomenon can manifest in different ways depending on the type of interaction network. In the mention network, users tend to engage more with others who share similar political biases, forming bias-based echo chambers. Meanwhile, in the retweet network, clustering occurs based on the reliability of the media sources that users amplify, leading to reliability-based echo chambers. These dynamics are analysed across key climate-related topics, including COP27, IPCC, Climate Refugees, and Doñana. Figures S11 and S12 illustrate these patterns, highlighting the structure and polarisation of online conversations.
@@
Figure S11: Detection of echo chambers in the mention network for political bias. Political bias of the neighbourhood as a function of the bias of a user in the mention network for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana.
@@
Figure S12: Detection of echo chambers in the retweet network for reliability. Media reliability of the neighbourhood as a function of the bias of a user in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. studied by considering only retweets.
Figure S13: Chamber overlap distribution $q_{ij}$. Chamber overlap distribution for (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana calculated on the top 2020 users of each community.
The chamber overlap can distinguish between users sharing similar echo chambers (Fig.S14). We constructed the user-to-user matrices where the values correspond to their chamber overlap. Using a hierarchical agglomerative algorithm, we clustered the distance matrix and obtained an almost perfect separation of users by communities. The separation between the users in the communities is neater in the COP27 and Doñana datasets compared to IPCC and Climate Refugees. Nevertheless, we observe a clear trend of similar echo chambers among users within the same community. In the COP27 network, the pattern is more evident for the users in the mainstream community (in blue). The separation between communities and the matching of echo chambers suggest a clearer separation in the Doñana dataset.
@@
Appendix C Policy and affective polarisation
S3 Hashtag analysis
Understanding the dynamics of hashtag usage across different communities provides valuable insights into the discourse surrounding key topics. In Fig. S15, we present the 20 most tweeted and retweeted hashtags for the two communities analyzed in each network. Our findings reveal a substantial overlap in hashtag usage between communities, with a dominance of generalistic hashtags. To further investigate these patterns, we assess the significance of differences in hashtag presence in Fig. S16. The results indicate that, in all cases, the variation in hashtag abundance between communities is statistically significant, highlighting distinct engagement patterns despite the observed overlap.